A balloon powered car is shot horizontally off the end of a 2.7 m high table at 5.5 m/s. How far horizontally from the edge of the desk does the car land?
1
Expert's answer
2020-12-21T10:39:11-0500

Time of falling "t" can be found from the kinematic equation (if there is no vertical initial velocity component):


"h = \dfrac{gt^2}{2}"

where "h = 2.7m" is the table height, and "g = 9.81m\/s^2" is the gravitational acceleration. Thus, obtain:



"t = \sqrt{\dfrac{2h}{g}} = \sqrt{\dfrac{2\cdot 2.7}{9.81}} \approx 0.74s"

In this time the survival kit covers the following horizontal distance:


"d = v_0t = 5.5m/s\cdot 0.74s \approx 4.1m"

Here "v_0 = 5.5m\/s" is the initial horizontal velocity of the car.


Answer. 4.1 m.